OpenStudy (anonymous):

Determine "m" such that the two vectors are orthogonal. 4mi=j, 9mi-25j

OpenStudy (anonymous):

(4mi+j).(9mi-25j)=0 (dot product) this gives u m=5/6

OpenStudy (anonymous):

and also m=-5/6

OpenStudy (anonymous):

Do you mind explaining how you did the problem? please.

OpenStudy (anonymous):

For two vectors to be orthogonal their dot product must be 0.

OpenStudy (anonymous):

Blah. dyslexic sorry.

OpenStudy (anonymous):

I understand it no that I can see it, there were no examples in my book. Thank you!

OpenStudy (anonymous):

\[<4m,1>\cdot <9m,-25> = 0\]\[\implies 36m^2 -25 = 0\]\[\implies 36m^2 = 25\]\[\implies m^2 = {25 \over 36} \implies m = \pm {5\over 6}\]

OpenStudy (anonymous):

There, that one is correct.